Compute the minimum distance that a truck slides on a horizontal

asphalt road if it is traveling at 20 m/s. The coefficient of sliding

friction between asphalt and rubber tire is 0.50. The mass of the

truck is 8000 kg.


1
Expert's answer
2022-03-28T11:02:26-0400

The truck acceleration

"a=\frac{F}{m}=\frac{\mu mg}{g}=\mu g\\
=0.50*9.8=4.9\:\rm m/s^2"

The distance

"d=\frac{v_0^2}{2a}=\frac{20^2}{2*4.9}=41\:\rm m"